# An ultra-fast intensity interferometry instrument

### v2

@@ 3.2 Bunching of A-type stars
Despite introducing an appropriate cable delay and taking appropriate measures to shield our electronics from the RF background of the telescope dome, we observed some deviation from shot noise when evaluating the RMSE for each of our measurements. The observed RMSE over the number of counts per bin $N$ is shown in Fig. 6. Ideally all RMSE curves should show shot noise closely following a square root curve. For Vega this curve is displayed in the top part of Fig. 6. Plotting the residuals divided by their shot noise value in the lower part of Fig. 6 shows that the relative residuals are initially close to zero for all observed stars, but around $N=0.25\times 10^{6}$ the deviations grow. Furthermore, the relative residuals are larger for targets at higher count rates, with Vega deviating 6% from shot noise at the end of its observation time, Altair deviating 1.5%, and Deneb nearly not deviating at all. While this discourages observations at higher detected photon rates with our current time tagger, we also note that the relative residuals seem to be rather stable once settled in, still allowing high S/N bunching measurements. Similar deviations from shot noise have been measured in our laboratory test (cf. Appendix B). We thus conclude that this deviation from shot noise is most likely intrinsic to our detection electronics.
@@
Figure 5: Bunching signal measured for three different stars shown in orange with corresponding error bars. The expected measurement result is shown as blue crosses. A Gaussian fit is used for centering purposes (green line). The measurement result fits well to the expected coherence time of 0.210.210.21\,0.21ps.
@@
Figure 6: Root mean square error (RMSE) (top) and re-scaled residuals of the RMSE from shot noise (bottom) plotted over the number of counts per bin $N$. The RMSE is expected to be shot noise $\mathrm{RMSE}=\sqrt{N}$ for $N$ counts per histogram bin. The upper plot shows deviations from shot noise in Vega’s curve for higher count rates. This is due to the limits of the detection system. These deviations are shown in greater detail for all observed stars in the lower plot. While the residuals are initially close to zero for all stars, the relative residuals are larger for targets with higher detected count rates. Even for the high S/N measurement recorded for Vega, the RMSE is at most 6% larger than the shot noise expectation.
Table 1: Overview of the observed A-type stars and their properties. Apparent magnitudes were taken from Bohlin & Gilliland (2004) and Ducati (2002).
@@ 3.3 Advantages of our setup
The short optical path of our setup, that is the distance from the entrance point of the light to the detector allows a straightforward alignment. The alignment is done by adjusting merely the telescope’s focus and its pointing/tracking of the star, in order for the beam to impinge centrally. All optical components were pre-aligned in the lab and no further adjustments of optics inside the setup were necessary.
The use of HPDs produced by Becker&Hickl has the advantage of observing stars in the B-band. This can contribute to new science as observations in the blue, here at 405405405\,405nm, are rather rare and can be used to test or verify commonly used models for stars in the blue (Sackrider et al., 2022) (Acharyya et al., 2024).
@@
Table 2 shows the observed count rates for all stars evaluated over all measurement nights. The median, maximum, and minimum count rate are calculated in a standard fashion, taking all data into account uncorrected. Note that the rather small minima for the count rates are due to bad weather conditions, these are. clouds covering the star for a short time during the night. The deviation in Table 2 is calculated after correcting for the deterministic change in count rate due to the star’s motion along the night sky. For all observed stars, the standard deviation of the count rate is orders of magnitude smaller than the median observed count rate, indicating fairly stable coupling of the star light to our detectors. For Vega and Deneb, we observe a larger fluctuation in count rate than for Altair, indicating slightly worse observation conditions. This is also reflected by the more pronounced loss in count rate when comparing minimum to median count rate for Vega and Deneb with respect to Altair.
Figure 7: Count rate at the detectors over time during the night of the 13th of August. The count rate is averaged every 303030\,30s. Three different stars were observed. One can clearly see that each measurement was quite stable, meaning. tracking of the star was never lost. This is due to the stability of the equatorial yoke mount of the telescope and the rather large active area of the HPDs. The time-evolutions of the count rates for Altair and Deneb are mainly driven by the air mass evolution due to sidereal motion. For Vega, the flux variation is more likely to be due to atmospheric transmission fluctuations.
@@ 4 Conclusion and outlook
A promising avenue for spatial photon correlation measurements is to duplicate our optical setup, replace the HPDs with Hi-QE based detectors and use a self-developed TDC. When performing a measurement for more than 101010\,10h using the quTAG, we saw that the re-scaled residuals increase compared to measurements at the telescope, see Appendix B and Fig. 11. This increase is believed to be due to the read-out electronics. As for measurements at the telescope we additionally saw that the residuals increase for higher count rates (Fig. 6). We are currently developing our own TDC aiming for a higher resolution at higher count rates. This TDC is dead time free (no additional 404040\,40ns) and therefore does not loose photon detection events so that the histogramming can be done via multi-start multi-stop. The new detection system would thus have an even lower timing jitter compared to the shown measurements.
@@
Data Availability
Acknowledgements.
We cordially thank the team of the Observatoire de la Côte d’Azur for letting us perform measurements at one of their telescopes. We also thank Nolan Matthews, Robin Kaiser, William Guerin and Mathilde Hugbart for helpful discussions on how to perform the measurement at the Omicron telescope. We acknowledge Oleg Kalekin for his support in recording the quantum efficiency measurements.
@@ Appendix A Timing jitter
In order to probe the timing resolution of our detection setup, we used laser pulses with a pulse duration much smaller than the expected timing resolution of the detection setup. We supply these pulses by single pass frequency tripling of a 156015601560\,1560nm IR pulsed laser (Toptica FemtoFErb 1560) with a nominal pulse duration $<80\,$fs using a PPLN crystal. Since the crystal is only 0.50.50.5\,0.5mm long, the tripled light at 520520520\,520nm has a nearly identical pulse duration as the fundamental pulses. After frequency tripling, the intensity of the laser is strongly attenuated, in order to ensure on average less than 1 photon per laser pulse being incident on each of the two detectors. With a 100100100\,100 MHz laser repetition rate the detectors would otherwise be saturated quickly. After attenuation the light is split and directed to the HPDs by the same 50:50 non-polarizing beam splitter used in the bunching measurements. At approx. 222\,2MHz count rate we observed a FWHM timing resolution of 42.442.442.4\,42.4ps. The result of this measurement is plotted in Fig. 8. This high resolution measurement is used to calculate the measurement expectation for the bunching measurements.
@@
Figure 8: Timing jitter measured with the full photo detection system of 2 HPDs, 2 CFDs and the quTAG, as well as one 50:50 non-polarizing beam splitter. The delta like pulses were supplied by a frequency tripled femtosecond IR laser. At approx. 222\,2MHz count rate a FWHM timing resolution of 42.442.442.4\,42.4ps is observed. The measurement is rescaled to have a unit integral.
@@ Appendix B Lab test results
In terms of measured coherence time, the laboratory test fits our expectations extremely well, with $(0.21\pm 0.02)\,$ps and an expected coherence time of 0.210.210.21\,0.21ps (cf. Fig. 9). To measure a normalized second order photon correlation function, such as the one shown in Fig. 9, some preprocessing is necessary. A raw correlation measurement is shown in Fig. 10. Note that here the second order correlation function $\Gamma^{(2)}(\tau_{\text{raw }})$ is not normalized. Besides the correlation peak (highlighted in green), some other regularly spaced peaks (highlighted in red) are visible. These peaks likely correspond to the frequency of the readout clock of the quTAG. The sections between these clock spikes are reasonably close to being shot noise limited. Thus, a 333\,3m cable delay is introduced such that the correlation peak lies inside one of these regions. The clock spikes shaded in red, as well as the region of the correlation peak shaded in green, were not used for the RMSE evaluations presented in this paper. Similarly, only the non-shaded ares of the raw correlation histogram were used in the determination of the average counts per bin, used to calculate the normalized second order correlation function.
Even in the laboratory, our detection electronics did not conform to shot noise perfectly. In fact, the deviations were more significant than at the telescope, finally saturating at a RMSE 15 per cent larger than the shot noise expectation (cf. Fig 11). The step in the relative RMSE at about $2\times 10^{6}$ counts per bin is likely due to the long uninterrupted uptime (more than 424242\,42h) of the detection system, as parameters such as the count rate or the ambient temperature did not fluctuate during the lab test.
Figure 9: Bunching signal measured in the laboratory using an XBO. The measurement result fits well to the expectation, giving a measured coherence time $\tau_{\mathrm{meas,2f-2f}}=(0.21\pm 0.02)\,$ps for an expected coherence time of 0.2120.2120.212\,0.212ps.
@@
Figure 10: Raw correlation histogram measured as a laboratory test. The area shaded in green is the area of the bunching peak, while the areas shaded in red signify the location of spikes of the quTAG’s readout clock. These areas were excluded in all RMSE evaluations.
@@
Figure 11: Root mean square error (RMSE) (top) and re-scaled residuals of the RMSE from shot noise (bottom) plotted over the number of coincidences per bin $N$. The RMSE is expected to be shot noise $\mathrm{RMSE}=\sqrt{N}$ for $N$ counts per histogram bin. The upper plot shows deviations from shot noise in a laboratory test for higher count rates. These deviations are shown in greater detail in the lower plot. Two nearly stable plateaus are visible. For smaller $N$ the residuals can be compared to the ones obtained at the telescope. However, due to the long measurement time in the laboratory test the residuals increase.

### v3

@@ 3.2 Bunching of A-type stars
Despite introducing an appropriate cable delay and taking appropriate measures to shield our electronics from the RF background of the telescope dome, we observed some deviation from shot noise when evaluating the RMSE for each of our measurements. The observed RMSE over the number of counts per bin $N$ is shown in Fig. 5(a). Ideally all RMSE curves should show shot noise closely following a square root curve. For Vega this curve is displayed in the top part of Fig. 5(a). Plotting the residuals divided by their shot noise value in the lower part of Fig. 5(a) shows that the relative residuals are initially close to zero for all observed stars, but around $N=0.25\times 10^{6}$ the deviations grow. Furthermore, the relative residuals are larger for targets at higher count rates, with Vega deviating 6% from shot noise at the end of its observation time, Altair deviating 1.5%, and Deneb nearly not deviating at all. While this discourages observations at higher detected photon rates with our current time tagger, we also note that the relative residuals seem to be rather stable once settled in, still allowing high S/N bunching measurements. Similar deviations from shot noise have been measured in our laboratory test (cf. Appendix B). We thus conclude that this deviation from shot noise is most likely intrinsic to our detection electronics.
@@
Figure 4: Bunching signal measured for different stars shown in orange with corresponding error bars. The expected measurement result is shown as blue crosses. A Gaussian fit is used for centering purposes (green line). The measurement result fits well to the expected coherence time of 0.210.210.21\,0.21ps.
@@
Figure 5: (a) Root mean square error (RMSE) (top) and re-scaled residuals of the RMSE from shot noise (bottom) plotted over the number of counts per bin $N$. The RMSE is expected to be shot noise $\mathrm{RMSE}=\sqrt{N}$ for $N$ counts per histogram bin. The upper plot shows deviations from shot noise in Vega’s curve for higher count rates. This is due to the limits of the detection system. These deviations are shown in greater detail for all observed stars in the lower plot. While the residuals are initially close to zero for all stars, the relative residuals are larger for targets with higher detected count rates. Even for the high S/N measurement recorded for Vega, the RMSE is at most 6% larger than the shot noise expectation. (b) Count rate at the detectors over time during the night of the 13th of August. The count rate is averaged every 303030\,30s. Three different stars were observed. One can clearly see that each measurement was quite stable, meaning. tracking of the star was never lost. This is due to the stability of the equatorial yoke mount of the telescope and the rather large active area of the HPDs. The time-evolutions of the count rates for Altair and Deneb are mainly driven by the air mass evolution due to sidereal motion. For Vega, the flux variation is more likely to be due to atmospheric transmission fluctuations.
Table 1: Overview of the observed A-type stars and their properties. Apparent magnitudes were taken from [24] and [25].
@@ 3.3 Advantages of our setup
The short optical path of our setup, that is the distance from the entrance point of the light to the detector allows a straightforward alignment. This is a major improvement compared to one of our previous setups that featured a deflection mirror at the telescope flange [17]. The alignment is done by adjusting merely the telescope’s focus and its pointing/tracking of the star, in order for the beam to impinge centrally. All optical components were pre-aligned in the lab and no further adjustments of optics inside the setup were necessary.
The use of HPDs produced by Becker&Hickl has the advantage of observing stars in the B-band. This can contribute to new science as observations in the blue, here at 405405405\,405nm, are rather rare and can be used to test or verify commonly used models for stars in the blue [26] [27].
@@
Table 2 shows the observed count rates for all stars evaluated over all measurement nights. The median, maximum, and minimum count rate are calculated in a standard fashion, taking all data into account uncorrected. Note that the rather small minima for the count rates are due to bad weather conditions, these are clouds covering the star for a short time during the night. The deviation in Table 2 is calculated after correcting for the deterministic change in count rate due to the star’s motion along the night sky. For all observed stars, the standard deviation of the count rate is orders of magnitude smaller than the median observed count rate, indicating fairly stable coupling of the star light to our detectors. For Vega and Deneb, we observe a larger fluctuation in count rate than for Altair, indicating slightly worse observation conditions. This is also reflected by the more pronounced loss in count rate when comparing minimum to median count rate for Vega and Deneb with respect to Altair.
@@ 4 Conclusion and outlook
A promising avenue for spatial photon correlation measurements is to duplicate our optical setup, replace the HPDs with Hi-QE based detectors and use a self-developed TDC. When performing a measurement for more than 101010\,10h using the quTAG, we saw that the re-scaled residuals increase compared to measurements at the telescope, see Appendix B and Fig. 7(b). This increase is believed to be due to the read-out electronics. As for measurements at the telescope we additionally saw that the residuals increase for higher count rates (Fig. 5(a)). We are currently developing our own TDC aiming for a higher resolution at higher count rates. This TDC is dead time free (no additional 404040\,40ns) and therefore does not loose photon detection events so that the histogramming can be done via multi-start multi-stop. The new detection system would thus have an even lower timing jitter compared to the shown measurements.
@@ Appendix A Timing jitter
In order to probe the timing resolution of our detection setup, we used laser pulses with a pulse duration much smaller than the expected timing resolution of the detection setup. We supply these pulses by single pass frequency tripling of a 156015601560\,1560nm IR pulsed laser (Toptica FemtoFErb 1560) with a nominal pulse duration $<80\,$fs using a PPLN crystal. Since the crystal is only 0.50.50.5\,0.5mm long, the tripled light at 520520520\,520nm has a nearly identical pulse duration as the fundamental pulses. After frequency tripling, the intensity of the laser is strongly attenuated, in order to ensure on average less than 1 photon per laser pulse being incident on each of the two detectors. With a 100100100\,100 MHz laser repetition rate the detectors would otherwise be saturated quickly. After attenuation the light is split and directed to the HPDs by the same 50:50 non-polarizing beam splitter used in the bunching measurements. At approx. 222\,2MHz count rate we observed a FWHM timing resolution of 42.442.442.4\,42.4ps. The result of this measurement is plotted in Fig. 6(a). This high resolution measurement is used to calculate the measurement expectation for the bunching measurements.
@@
Figure 6: (a) Timing jitter measured with the full photo detection system of 2 HPDs, 2 CFDs and the quTAG, as well as one 50:50 non-polarizing beam splitter. The delta like pulses were supplied by a frequency tripled femtosecond IR laser. At approx. 222\,2MHz count rate a FWHM timing resolution of 42.442.442.4\,42.4ps is observed. The measurement is rescaled to have a unit integral. (b) Bunching signal measured in the laboratory using an XBO. The measurement result fits well to the expectation, giving a measured coherence time $\tau_{\mathrm{meas,2f-2f}}=(0.21\pm 0.02)\,$ps for an expected coherence time of 0.2120.2120.212\,0.212ps.
@@ Appendix B Lab test results
In terms of measured coherence time, the laboratory test fits our expectations extremely well, with $(0.21\pm 0.02)\,$ps and an expected coherence time of 0.210.210.21\,0.21ps (cf. Fig. 6(b)). To measure a normalized second order photon correlation function, such as the one shown in Fig. 6(b), some preprocessing is necessary. A raw correlation measurement is shown in Fig. 7(a). Note that here the second order correlation function $\Gamma^{(2)}(\tau_{\text{raw }})$ is not normalized. Besides the correlation peak (highlighted in green), some other regularly spaced peaks (highlighted in red) are visible. These peaks likely correspond to the frequency of the readout clock of the quTAG. The sections between these clock spikes are reasonably close to being shot noise limited. Thus, a 333\,3m cable delay is introduced such that the correlation peak lies inside one of these regions. The clock spikes shaded in red, as well as the region of the correlation peak shaded in green, were not used for the RMSE evaluations presented in this paper. Similarly, only the non-shaded ares of the raw correlation histogram were used in the determination of the average counts per bin, used to calculate the normalized second order correlation function.
Even in the laboratory, our detection electronics did not conform to shot noise perfectly. In fact, the deviations were more significant than at the telescope, finally saturating at a RMSE 15 percent larger than the shot noise expectation (cf. Fig 7(b)). The step in the relative RMSE at about $2\times 10^{6}$ counts per bin is likely due to the long uninterrupted uptime (more than 424242\,42h) of the detection system, as parameters such as the count rate or the ambient temperature did not fluctuate during the lab test.
@@
Figure 7: (a) Raw correlation histogram measured as a laboratory test. The area shaded in green is the area of the bunching peak, while the areas shaded in red signify the location of spikes of the quTAG’s readout clock. These areas were excluded in all RMSE evaluations. (b) Root mean square error (RMSE) (top) and re-scaled residuals of the RMSE from shot noise (bottom) plotted over the number of coincidences per bin $N$. The RMSE is expected to be shot noise $\mathrm{RMSE}=\sqrt{N}$ for $N$ counts per histogram bin. The upper plot shows deviations from shot noise in a laboratory test for higher count rates. These deviations are shown in greater detail in the lower plot. Two nearly stable plateaus are visible. For smaller $N$ the residuals can be compared to the ones obtained at the telescope. However, due to the long measurement time in the laboratory test the residuals increase.
@@ Appendix C HPD quantum efficiency
Disclosures
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.
Code, Data, and Materials Availability
Acknowledgments
We cordially thank the team of the Observatoire de la Côte d’Azur for letting us perform measurements at one of their telescopes. We also thank Nolan Matthews, Robin Kaiser, William Guerin and Mathilde Hugbart for helpful discussions on how to perform the measurement at the Omicron telescope. We acknowledge Oleg Kalekin for his support in recording the quantum efficiency measurements. We acknowledge the financial support of the French National Research Agency (project I2C, ANR-20-15CE31-0003).
Biographies and photographs of the other authors are not available.